OpenStudy (anonymous):

Which of the following points is a solution to the following system of equations? x + y = 1 3x – y = 3 A. (3, 0) B. (5, –1) C. (1, 4) D. (1, 0)

OpenStudy (timaashorty):

A. 3 + 1 = 4 nope (so its not the first one obvously) B. 5 + (-1) = 4 (not this one) C. 1 + 4 = 5 ( not this one) \[\ \huge D. 1+0 = 1 ; 3 (1) - (0) = 3 \ This\ one. \]

OpenStudy (timaashorty):

all you needed to do is plug in the answer choices.